José Antonio de la Peña: On the inductive construction of Galois coverings of algebras

Let A be a finite dimensional algebra over an algebraically closed field k. Assume that A is given as a one-point extension A = B[M], for an algebra B and a B-module M. We consider the general problem of construction and properties of coverings of A knowing the coverings of B. In particular, we use this technique to prove that the fundamental group of certain tame algebras is free.